f(x)=x+2/x-6 i
is the point on (3,14)on the graph of f

Answer by jim_thompson5910(35256)

f(x)=(x+2)/(x-6)f(3)=(3+2)/(3-6)

f(3)=5/(-3)

f(3)=-5/3

So the point (3,14) is NOT on the graph of f.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------